# Unsupervised Cross-Domain Image Retrieval via Prototypical Optimal Transport

### v1

@@ Abstract
Unsupervised cross-domain image retrieval (UCIR) aims to retrieve images sharing the same category across diverse domains without relying on labeled data. Prior approaches have typically decomposed the UCIR problem into two distinct tasks: intra-domain representation learning and cross-domain feature alignment. However, these segregated strategies overlook the potential synergies between these tasks. This paper introduces ProtoOT, a novel Optimal Transport formulation explicitly tailored for UCIR, which integrates intra-domain feature representation learning and cross-domain alignment into a unified framework. ProtoOT leverages the strengths of the K-means clustering method to effectively manage distribution imbalances inherent in UCIR. By utilizing K-means for generating initial prototypes and approximating class marginal distributions, we modify the constraints in Optimal Transport accordingly, significantly enhancing its performance in UCIR scenarios. Furthermore, we incorporate contrastive learning into the ProtoOT framework to further improve representation learning. This encourages local semantic consistency among features with similar semantics, while also explicitly enforcing separation between features and unmatched prototypes, thereby enhancing global discriminativeness. ProtoOT surpasses existing state-of-the-art methods by a notable margin across benchmark datasets. Notably, on DomainNet, ProtoOT achieves an average P@200 enhancement of 24.44%, and on Office-Home, it demonstrates a P@15 improvement of 12.12%. Code is available at https://github.com/HCVLAB/ProtoOT.
@@ Introduction
  • ProtoOT outperforms existing state-of-the-art techniques by a substantial margin across benchmark datasets. Notably, on DomainNet, ProtoOT achieves an average enhancement of 24.44% in terms of P@200, and on Office-Home, an improvement of 12.12% in terms of P@15.
@@
Unsupervised domain adaptation (UDA) aims to address the challenge of transferring knowledge from a labeled source domain to a target domain without any annotations. To mitigate the domain shift, existing methods focus on reducing domain discrepancy through techniques like Maximum Mean Discrepancy (Gretton et al. 2012; Long et al. 2017), domain adversarial training (Zhang et al. 2018; Dai et al. 2020) and prototypes alignment(Lin et al. 2022; Li, Lü, and Li 2022; Yue et al. 2021; Chang et al. 2022). However, while these methods have demonstrated substantial efficacy in UDA, they cannot be used directly for UCIR tasks because there are no labels in the source domain. In this paper, we delve into a more challenging scenario in which both the source and target domains are unlabeled. The problem of exploring weak self-representation learning for both domains while concurrently transferring knowledge from the other remains unresolved.
@@
Optimal transport (OT), initially proposed by Kantorovich (Villani et al. 2009), offers an efficient solution for transferring mass from one distribution to another. In the field of computer vision, OT has been utilized in diverse tasks like point cloud registration(Shen et al. 2021), shape matching(Saleh et al. 2022) and learning with noisy labels(Chang, Shi, and Wang 2023). These tasks involve seeking an optimal mapping between data distributions to minimize a cost function. Notably, (Asano, Rupprecht, and Vedaldi 2019; Caron et al. 2020) have achieved significant advancements in unsupervised representation learning through OT-driven matching of samples to prototypes. Furthermore, in the context of domain adaptation, OT-based approaches(Courty, Flamary, and Tuia 2014; Courty et al. 2017; Redko et al. 2019; Chang et al. 2022) have exhibited promise in mitigating the gap between source and target domains by aligning their respective probability distributions. However, jointly exploring self-supervised feature representation learning and domain adaptation within a unified OT framework remains unexplored.
@@
[…]
@@
[…]
@@
where the symbol ${}^{\prime}$ denotes a different domain, $\bm{\beta}_{cr}$ is the categories distribution of $\bm{M}$ obtained by K-means. We consider the matched prototype $\bm{C}_{y^{\prime}_{i}}$ as the positive sample and other prototypes are treated as negative samples $\bm{n}^{\prime}\in\mathcal{N}^{cr}_{i}$.
@@ A Unified ProtoOT Framework for UCIR
P@50 37.07 37.21 36.89 41.23 47.09 68.68 +21.59
DomainNet P@100 33.34 32.59 31.84 36.87 43.47 67.04 +23.57
P@200 28.72 27.85 26.69 31.74 39.09 63.53 +24.44
P@1 33.18 35.49 37.17 38.07 40.67 48.61 +7.94
Office-Home P@5 29.76 32.15 34.63 35.01 37.60 47.88 +10.28
P@15 25.78 28.30 31.30 31.42 34.15 46.27 +12.12
@@
We employ the ResNet-50(He et al. 2016) architecture as the encoder $f_{\theta}$, initializing its parameters using MoCov2(Chen et al. 2020) trained on the unlabeled ImageNet dataset(Deng et al. 2009). The extracted features undergo $l_{2}$-normalization. Our optimization employs the Adam optimizer with a learning rate of $2.5\times 10^{-4}$ over 200 epochs, with a batch size of 64. The K-means clustering is performed on the memory bank at the end of each epoch. For the Sinkhorn Algorithm(Cuturi 2013), the entropic regularization coefficient $\epsilon$ is set to 0.05 and following (Caron et al. 2020) the iterations is 3. The number of prototypes corresponds to the number of classes in the training set: 65 for Office-Home and 7 for DomainNet. Implementation of our framework is in PyTorch(Paszke et al. 2019).
@@
ProtoOT’s impressive performance is evident from the summarized results in Table 3. On the Office-Home dataset, ProtoOT outperforms the leading baseline(Hu and Lee 2022) with average improvements of 7.94%, 10.28%, and 12.12% in P@1, P@5, and P@15, respectively. The improvements are even more pronounced on DomainNet, showcasing substantial enhancements of 21.59%, 23.57%, and 24.44% in P@50, P@100, and P@200. Table 1 and 2 show the detailed retrieval performance for Office-Home and DomainNet, respectively. It is worth noting that the Painting $\rightarrow$ Clipart retrieval on DomainNet shows the most significant improvement, with an accuracy boost of 39.95% in terms of P@200. Furthermore, ProtoOT maintains over 50% retrieval precision consistently, even in challenging tasks that involve Quickdraw, with an impressive average enhancement of 22.62%, 21.74%, and 19.90% in P@50, P@100, and P@200, respectively. This is attributed to the strong matching capabilities of ProtoOT.
@@
SOT 54.92 51.57 46.13
SOT SOT 62.23 59.91 56.90
SOT ProtoOT 67.56 65.80 61.97
ProtoOT 55.22 52.57 47.68
ProtoOT SOT 66.86 65.29 61.84
ProtoOT ProtoOT 68.68 67.04 63.53
@@
To further validate the superiority of our ProtoOT, we conducted a comparison with several common variants of OT. The Unbalanced OT(Séjourné, Peyré, and Vialard 2023) uses soft penalties to relax the conservation of marginal constraints. The Partial OT(Figalli 2010) focuses on transporting only a fraction of mass at the lowest cost. Due to their intrinsic characteristics that are not well suited for the UCIR task, as depicted in Table 5, they yield suboptimal performance.

### v4

@@ Abstract
Unsupervised cross-domain image retrieval (UCIR) aims to retrieve images sharing the same category across diverse domains without relying on labeled data. Prior approaches have typically decomposed the UCIR problem into two distinct tasks: intra-domain representation learning and cross-domain feature alignment. However, these segregated strategies overlook the potential synergies between these tasks. This paper introduces ProtoOT, a novel Optimal Transport formulation explicitly tailored for UCIR, which integrates intra-domain feature representation learning and cross-domain alignment into a unified framework. ProtoOT leverages the strengths of the K-means clustering method to effectively manage distribution imbalances inherent in UCIR. By utilizing K-means for generating initial prototypes and approximating class marginal distributions, we modify the constraints in Optimal Transport accordingly, significantly enhancing its performance in UCIR scenarios. Furthermore, we incorporate contrastive learning into the ProtoOT framework to further improve representation learning. This encourages local semantic consistency among features with similar semantics, while also explicitly enforcing separation between features and unmatched prototypes, thereby enhancing global discriminativeness. ProtoOT surpasses existing state-of-the-art methods by a notable margin across benchmark datasets. Notably, on DomainNet, ProtoOT achieves an average P@200 enhancement of 18.17%, and on Office-Home, it demonstrates a P@15 improvement of 3.83%.
@@ Introduction
  • ProtoOT outperforms existing state-of-the-art techniques by a substantial margin across benchmark datasets. Notably, on DomainNet, ProtoOT achieves an average enhancement of 18.17% in terms of P@200, and on Office-Home, an improvement of 3.83% in terms of P@15.
@@
Unsupervised domain adaptation (UDA) aims to address the challenge of transferring knowledge from a labeled source domain to a target domain without any annotations. To mitigate the domain shift, most existing methods have focused on reducing domain discrepancy through techniques like Maximum Mean Discrepancy (Gretton et al. 2012; Long et al. 2017), domain adversarial training (Zhang et al. 2018; Dai et al. 2020) and prototypes alignment(Lin et al. 2022; Li, Lü, and Li 2022; Yue et al. 2021; Chang et al. 2022). However, while these methods have demonstrated substantial efficacy in UDA, they cannot be used directly for UCIR tasks because there are no labels in the source domain. In this paper, we delve into a more challenging scenario in which both the source and target domains are unlabeled. The problem of exploring weak self-representation learning for both domains while concurrently transferring knowledge from the other remains unresolved.
@@
Optimal transport (OT), initially proposed by Kantorovich (Villani et al. 2009), offers an efficient solution for transferring mass from one distribution to another. The development of optimized solvers(Cuturi 2013; Lahn, Raghvendra, and Zhang 2024) have provided a guarantee for the wide application of OT. In the field of computer vision, OT has been utilized in diverse tasks like point cloud registration(Shen et al. 2021) and learning with noisy labels(Chang, Shi, and Wang 2023). These tasks involve seeking an optimal mapping between data distributions to minimize a cost function. Notably, (Asano, Rupprecht, and Vedaldi 2019; Caron et al. 2020) have achieved significant advancements in unsupervised representation learning through OT-driven matching of samples to prototypes. Furthermore, in the context of domain adaptation, OT-based approaches(Courty, Flamary, and Tuia 2014; Courty et al. 2017; Redko et al. 2019; Chang et al. 2022) have exhibited promise in mitigating the gap between source and target domains by aligning their respective probability distributions. However, jointly exploring self-supervised feature representation learning and domain adaptation within a unified OT framework remains unexplored.
@@
where $\varepsilon>0$, the entropic regularization term $\textit{H}(\bm{Q})=-\sum_{ij}\bm{Q}_{ij}\log\bm{Q}_{ij}$.
@@
$\bm{Q}^{*}(\bm{M},\bm{C},\bm{\beta})=\operatorname*{argmax}_{\bm{Q}\in\bm{U}(\bm{\beta})}\text{Tr}(\bm{Q}^{\top}\bm{M}\bm{C}^{T})+\varepsilon\textit{H}(\bm{Q}),$ (3)
@@
where the symbol denotes a different domain, $\bm{\beta}_{cr}$ is the categories distribution of $\bm{M}$ obtained by K-means. We consider the matched prototype $\bm{C}_{y^{\prime}_{i}}$ as the positive sample and other prototypes are treated as negative samples $\bm{n}^{\prime}\in\mathcal{N}^{cr}_{i}$.
@@ A Unified ProtoOT Framework for UCIR
P@50 37.07 37.21 36.89 41.23 47.09 64.37 +17.28
DomainNet P@100 33.34 32.59 31.84 36.87 43.47 61.93 +18.46
P@200 28.72 27.85 26.69 31.74 39.09 57.26 +18.17
P@1 33.18 35.49 37.17 38.07 40.67 43.85 +3.18
Office-Home P@5 29.76 32.15 34.63 35.01 37.60 41.29 +3.69
P@15 25.78 28.30 31.30 31.42 34.15 37.98 +3.83
@@
We employ the ResNet-50(He et al. 2016) architecture as the encoder $f_{\theta}$, initializing its parameters using MoCov2(Chen et al. 2020) trained on the unlabeled ImageNet dataset(Deng et al. 2009). The extracted features undergo $l_{2}$-normalization. Our optimization employs the Adam optimizer with a learning rate of $2.5\times 10^{-4}$ over 200 epochs, with a batch size of 64. The K-means clustering is performed on the memory bank at the end of each epoch. For the Sinkhorn Algorithm(Cuturi 2013), the entropic regularization coefficient $\epsilon$ is set to 0.05 and following (Caron et al. 2020) the iterations is 3. The number of prototypes corresponds to the number of classes in the training set: 65 for Office-Home and 7 for DomainNet. In the initial phase of training, we employ the same loss function as used in MoCov2(Chen et al. 2020) for per-domain pre-training.
@@
ProtoOT’s impressive performance is evident from the summarized results in Table 3. On the Office-Home dataset, ProtoOT outperforms the leading baseline(Hu and Lee 2022) with average improvements of 3.18%, 3.69%, and 3.83% in P@1, P@5, and P@15, respectively. The improvements are even more pronounced on DomainNet, showcasing substantial enhancements of 17.28%, 18.46%, and 18.17% in P@50, P@100, and P@200. Table 1 and 2 show the detailed retrieval performance for Office-Home and DomainNet, respectively. It is worth noting that the Painting $\rightarrow$ Clipart retrieval on DomainNet shows the most significant improvement, with an accuracy boost of 30.54% in terms of P@100. Furthermore, ProtoOT maintains over 50% retrieval precision consistently, even in challenging tasks that involve Quickdraw, with an impressive average enhancement of 20.40%, 20.51%, and 20.47% in P@50, P@100, and P@200, respectively. This is attributed to the strong matching capabilities of ProtoOT.
@@
SOT 55.72 52.41 46.58
SOT SOT 60.08 57.01 51.46
SOT ProtoOT 60.58 57.78 52.16
ProtoOT 56.51 53.13 47.02
ProtoOT SOT 62.83 60.29 55.43
ProtoOT ProtoOT 64.37 61.93 57.26
@@
To further validate the superiority of our ProtoOT, we conducted a comparison with several common variants of OT. The Unbalanced OT(Séjourné, Peyré, and Vialard 2023) uses soft penalties to relax the conservation of marginal constraints. The Partial OT(Figalli 2010; Phatak et al. 2022) focuses on transporting only a fraction of mass at the lowest cost. Due to their intrinsic characteristics that are not well suited for the UCIR task, as depicted in Table 5, they yield suboptimal performance.
@@ Acknowledgments
Supplementary materials
Pre-training Details
In the initial phase of training, we employ the same loss function as used in MoCov2(Chen et al. 2020) for per-domain pre-training. The loss of the pre-training stage is as follow:
$\mathcal{L}_{pre}=\sum\limits_{i\in\mathcal{D}}-\log\frac{\exp(\bm{q}^{T}_{i}\bm{q}^{\prime}_{i}/\tau)}{\sum\limits_{\bm{j}\in\mathcal{D}}\exp(\bm{q}^{T}_{i}\bm{q}^{\prime}_{j}/\tau)},$ (13)
where $\bm{q}_{i}$ and $\bm{q}^{\prime}_{i}$ can be the image features from different augmented views of image $i$, $\tau$ is the temperature value. Specifically, during the initial 30 epochs of training, the pre-training loss term is incorporated to facilitate feature learning. Beyond the 30th epoch, this instance-wise loss is omitted to promote the acquisition of the learned features with more distinct class-level semantic separability.
Visualization
To provide a more intuitive comparison between our ProtoOT and DD(Hu and Lee 2022), we visualize Clipart-Sketch and Painting-Clipart on DomainNet. As shown in Figgure 4, we can observe that compared to DD, our ProtoOT yields more compact intra-class features and distinct inter-class separations.This can eliminate interference from other categories, thus contributing to the effectiveness of our ProtoOT approach. In Figure 5 and 6, we present additional retrieval results to demonstrate the effectiveness of our ProtoOT.
(a) DD (Clipart-Sketch)
(b) ProtoOT (Clipart-Sketch)
(c) DD (Painting-Clipart)
(d) ProtoOT (Painting-Clipart)
Figure 4: 2-d t-SNE visualizations feature representations learned by DD and our proposed ProtoOT on DomainNet. Each class is represented by a number and Each sample is colored by its corresponding domain.
Figure 5: Top 15 retrieval results in DomainNet. The green and red boxes denote correct and incorrect retrievals, respectively.
Figure 6: Top 15 retrieval results in Office-Home. The green and red boxes denote correct and incorrect retrievals, respectively.